# Symmetric square type L–series

### v1

@@ 1. Introduction
Symmetric square $L$-series have already been constructed by Shimura [ShimuraMF_HalfIntegralWeight]. These were instrumental in establishing the Shimura lift – a pivotal correspondence between modular forms of half-integral and integral weights. This groundbreaking work catalysed further investigations, including Shimura’s own extensions [Shimura1975] and generalisations to higher rank groups in [Bump1992]. The present work addresses the natural extension of these $L$-series to the setting of vector-valued modular forms associated to the Weil representation. More explicitly, let $(L,\operatorname{q})$ denote a non-degenerate even lattice, $L^{\prime}$ denote its dual and $k\in\tfrac{1}{2}\mathbb{Z}$ a half-integral number. Further, if $\mathcal{M}_{L,k}$ represents the space of $\mathbb{C}[L^{\prime}/L]$ valued holomorphic modular forms of weight $k$ for the associated Weil representation, then any form $f\in\mathcal{M}_{L,k}$ has a Fourier expansion
@@
In extending symmetric square $L$-series from the scalar to the vector-valued setting, one might initially consider associating an $L$-series with each component function of $f$ individually, meaning for fixed $\mu\in L^{\prime}/L$ in (1.1). However, closer examination reveals that the natural analogues exhibit intertwining of the components. For clarity of exposition, we present a special instance of the series discussed in the main body of the paper. For an anisotropic vector $\ell\in L^{\prime}$, where $(L,\operatorname{q})$ denotes the base lattice, we define
(1.2) $L_{\ell}(f,s)=\sum_{n\in\mathbb{N}}\frac{a(n\overline{\ell},n^{2}\operatorname{q}(\ell))}{n^{s}}$
@@
A key feature of Shimura’s $L$-series is their admission of infinite product expansions when they are attached to Hecke eigenforms. We replicate this property in the vector-valued setting through careful analysis of Hecke operators, building upon the work of Bruinier and Stein (cf. [BruinierStein], [SteinHecke]). In the process, a bound on eigenvalues is derived (cf. Lemma 4.11). We state a simplified version of Theorem 5.12.
@@
=\;= $a(\lambda,\operatorname{q}(\ell))\cdot\prod_{p\nmid\operatorname{lev}(L)}\frac{1+\delta_{p\nmid\operatorname{q}(\ell)}\frac{G_{\mathcal{L}}(1)}{G_{\mathcal{L}}(p)}\cdot p^{k-1-s}}{1-\left(\frac{\sigma_{p}}{p^{k-1}}-(1-p^{-1})\frac{G_{\mathcal{L}}(1)}{G_{\mathcal{L}}(p)}\right)\cdot p^{k-1-s}+p^{2(k-1-s)}}$
@@ 2. Setting
We let $(L,\operatorname{q})$ denote a non-degenerate even quadratic lattice of signature $({m^{+}},{m^{-}})$ and denote its bilinear form by $\operatorname{b}$ such that $\operatorname{q}(x)\coloneq\operatorname{b}(x,x)/2$. The dual lattice $L^{\prime}\subset L\otimes_{\mathbb{Z}}\mathbb{Q}$ represents $\operatorname{Hom}_{\mathbb{Z}}(L,\mathbb{Z})$ via $\operatorname{b}$ and we write $\operatorname{lev}(L)$ for the level of the lattice, being the smallest positive integer such $(L^{\prime},\operatorname{lev}(L)\cdot\operatorname{q})$ is even. Recall the inclusion $L\leq L^{\prime}$ and note that :-$\mathcal{L}\coloneq L^{\prime}/L$ is a finite abelian group whose order is the absolute value of a representing matrix of $\operatorname{b}$. Further, the quadratic form $\operatorname{q}$ descends to a quadratic form $\operatorname{\overline{q}}:\mathcal{L}\to\mathbb{Q}/\mathbb{Z}$ that is non-degenerate and such a pair $(\mathcal{L},\operatorname{\overline{q}})$ is called a discriminant form. In fact, any discriminant form arises from a non-degenerate even lattice in the above fashion. If we set $\operatorname{sig}(\mathcal{L})={m^{+}}-{m^{-}}\mod 8$ for some lattice $L$ inducing $\mathcal{L}$, this quantity is well defined by Milgram’s formula. For a natural number $n\in\mathbb{N}$ denote by $L(n)$ the pair $(L,n\cdot\operatorname{q})$ and set $\mathcal{L}(n)\coloneq L(n)^{\prime}/L(n)$.
@@
(2.1) 00{0}${{{}_{n}\mathcal{L}}}$${\mathcal{L}}$${{{}^{n}\!\mathcal{L}}}$00{0}$\scriptstyle{\iota}$$\scriptstyle{\cdot n}$
and find that ${{}^{n}\!\mathcal{L}}$ is exactly the orthogonal complement of ${{}_{n}\mathcal{L}}$. We denote by $\mathbb{C}[\mathcal{L}]$ the group algebra of $\mathcal{L}$ with standard basis $\mathfrak{e}_{\lambda}$ for $\lambda\in\mathcal{L}$. There is a representation of $\operatorname{SL}_{2}(\mathbb{Z})$ on the group algebra $\mathbb{C}[\mathcal{L}]$ first described by Schoeneberg. In fact, to obtain a well-defined representation when $\operatorname{sig}(\mathcal{L})$ is odd, it must be considered on the metaplectic extension, rather than on $\operatorname{SL}_{2}$ directly. Let $\R\operatorname{\overline{GL_{2}^{+}}}(\R)$ denote the metaplectic double cover of $\operatorname{GL}_{2}(\R)^{+}$. A model is given by pairs $(\gamma,\phi)$, where $\gamma\in\operatorname{GL}_{2}(\R)^{+}$ and $\phi:\mathbb{H}\to\mathbb{C}$ is a holomorphic square root of the factor of automorphy $j(\gamma,\tau)$, meaning $\phi^{2}(\tau)=j(\gamma,\tau)$. Multiplication is then declared by
@@
[…]
@@
In case of odd rank, the oddity formula [ConwaySloane1998, Chap. 15 p. 383 (30)] implies $4\mid N$, in particular, $\mathcal{L}$ contains $2$-adic Jordan components. In this case, there is a well known section
@@
with $\epsilon_{d}=1$ or $i$, depending on whether $d\equiv 1$ or $3\mod 4$. The argument at the end of [Borcherds2000, Thm. 5.4 p. 330] yields that ${\rho}_{\mathcal{L}}$ is trivial on $s(\Gamma(N))$ and factors through the central extension of $\operatorname{SL}_{2}(\mathbb{Z}/N\mathbb{Z})$ by {±$\{\pm 1\}$ given by
@@
In fact, the section in (2.5) may be considered on $\Gamma_{0}(4)$, provided a larger central extension of $\R\operatorname{SL}_{2}(\R)$ is considered which is given in [ShimuraMF_HalfIntegralWeight, p. 444] by allowing an additional phase for the root of the factor of automorphy and will be denoted by $\mathcal{G}$.
@@
Let $\R\gamma\in\operatorname{GL}_{2}^{+}(\R)$ and $f:\mathbb{H}\to\mathbb{C}$ be a function. Define
@@
Finally, a holomorphic modular form $f:\mathbb{H}\to\mathbb{C}$ of weight $k$ for some group $\Gamma\leq\Gamma(1)$ with character $\chi:\Gamma\to\mathbb{T}$ is a holomorphic function such that for all $\gamma\in\Gamma$ we have $f|_{k}\gamma=\chi(\gamma)f$ and $f$ is holomorphic at the cusps of $\Gamma$. The space of such functions is denoted by $\mathcal{M}_{k}(\Gamma,\chi)$ and the subspace of elements vanishing at the cusps is denoted by $\mathcal{S}_{k}(\Gamma,\chi)$ and called the space of cusp forms. For the technical modifications required for forms of half-integral weight, we refer to [ShimuraMF_HalfIntegralWeight, p. 444]. Note that Definition 2.3 may also be generalised to the extension $\mathcal{G}$ of $\R\operatorname{GL}_{2}^{+}(\R)$ that Shimura uses.
@@ 2.1. Gauss sums
This is a complementary section containing different Gauss sums appearing in section 4. The presented sums and some relations are already found in [SteinHecke, Sec. 4] including further computations which we will not repeat.
@@
Let $(L,\operatorname{q})$ be a quadratic $\mathbb{Z}$-lattice of rank $m\in\mathbb{N}$. For aprime $p$, consider $h\in\mathbb{Z}$ being coprime to $p$ and $n,d\in\mathbb{N}$ plus a Dirichlet Character $\chi$ modulo $n$. Define
@@
(2.9) $\epsilon_{d}=\begin{cases}1,&\textnormal{if }d\equiv 1\mod 4,\\ i,&\textnormal{if }d\equiv 3\mod 4.\\ \end{cases}$
@@
$g_{p}\left[n,\chi_{1},h\right]$ =$=\begin{cases}p^{n-1}(p-1),&\textnormal{ if }\nu_{p}(h)\geq n,\\ -p^{n-1},&\textnormal{ if }\nu_{p}(h)=n-1,\\ 0,&\textnormal{ if }\nu_{p}(h)<n-1.\end{cases}$
$g_{p}\left[n,\chi_{p},h\right]$ =$=\begin{cases}p^{n-1/2}\left(\frac{h/p^{n-1}}{p}\right)\epsilon_{p},&\textnormal{ if }\nu_{p}(h)=n-1,\\ 0,&\textnormal{ if }\nu_{p}(h)\neq n-1.\end{cases}$
@@
Let $(L,\operatorname{q})$ be an integral $\mathbb{Z}$-lattice. By [SteinHecke, Lemma 4.3] for any $n\in\mathbb{N}$ and odd prime $p$, there is an orthogonal decomposition of $L/p^{n}L$ into $\mathbb{Z}/p^{n}\mathbb{Z}$ submodules
@@
where $L_{i}=(\mathbb{Z}/p^{n}\mathbb{Z})v_{i}$ is one dimensional with $\operatorname{b}(v_{i},v_{i})\in(\mathbb{Z}/p^{n}\mathbb{Z})^{\times}$, $M_{j}=(\mathbb{Z}/p^{n}\mathbb{Z})v_{j}$ with $\operatorname{b}(v_{j},v_{j})\in p^{k}(\mathbb{Z}/p^{n}\mathbb{Z})^{\times}$ where $11\leq k\leq n-$ 1, and $\operatorname{b}(N,N)\subseteq p^{n}\mathbb{Z}$. We may assume the $M_{j}$ being sorted with respect to increasing valuation $k$ of $\operatorname{q}(v_{j})$. Define
@@
In [BruinierStein], the authors compare [McGraw2003, Lem. 4.6 p. 115] to [Borcherds2000, Thm. 5.4 p. 329] in order to conclude that for $\gcd(d,\operatorname{lev}(\mathcal{L}))=1$ the following association defines a character:
@@
Let $\mathcal{L}$ be the discriminant form of a lattice $L$ of rank $m$ and $d\in\mathbb{N}$. Then, by [Barnard2003, Thm. 5.2.2] we find
@@
  • Proof:
    Only the last equation has to be proven. By [ScheithauerWeilRep, Prop. 3.8] we find the identity $|G_{\mathcal{L}}(d)|=\sqrt{|{{}_{d}\mathcal{L}}||\mathcal{L}|}$. As a consequence,
@@
Do note that explicit formulae for a more general version of $G_{\mathcal{L}}$ are also contained in [ScheithauerWeilRep, Thm. 3.9 p. 11].
@@ 3. Modular forms
where $\overline{\mu}$ refers to the image of $\mu\in L^{\prime}/M$ under the projection to $L^{\prime}/L$ (cf. [Scheithauer2015, Sec. 4]). This construction plays a role in Subsection 5.1 and there are further means to construct induce vector-valued modular forms from existing ones described in [Scheithauer2015, Sec. 4]. However, a concrete class which is vital for our purposes is the following.
@@
(3.3) $\theta_{L,\lambda}(\tau)\coloneq\;$ :- $v^{\tfrac{{m^{-}}}{2}}\cdot\sum_{l\in L+\lambda}e\left(\tau\operatorname{q}(l)\right),\qquad\Theta_{L}\coloneq\sum_{\lambda\in L^{\prime}/L}\theta_{L,\lambda}(\tau)\mathfrak{e}_{\lambda}.$
These theta series $\Theta_{L}$ are modular forms of weight ${m^{+}}/2$ (cf. [Borcherds1998, Thm. 4.1 p. 505]) and their Fourier coefficients equal representation numbers of the lattice $L$. Comparing these theta forms to Eisenstein series will yield formulae for the representation numbers in terms of divisor functions up to the coefficients of a cusp form. Thus, deriving asymptotic bounds on the Fourier coefficients of cusp forms is vital for quantifying these formulae. Additionally, these bounds serve, in our context, as a basis for proving convergence results for $L$-series associated to cusp forms. We will reduce this question to the scalar-valued case and begin with stating the following observation.
Proposition 3.3 ([Borcherds2000, Thm. 5.4]).
@@
where $\chi_{\mathcal{L}}$ denotes the character of $\overline{\Gamma}_{0}(N)$ defined in [Borcherds2000, Thm. 5.4].
In fact, in case of $4\nmid N$, the character $\chi_{\mathcal{L}}$ is trivial on the preimage of $\Gamma(N)$ in $\operatorname{Mp}_{2}(\mathbb{Z})$. In the other case, meaning $4\mid N$, the section $s(\Gamma(N))$ in $\operatorname{Mp}_{2}(\mathbb{Z})$ is contained in the kernel of the character (cf. [Borcherds2000, Lem. 5.3 p. 329]). Hence, the component functions $f_{\lambda}$ of a vector-valued modular form $f\in\mathcal{M}_{L,k}$ are actually contained in $\mathcal{M}_{k}(\Gamma(\operatorname{lev}(L)))$.
As a consequence, for the purpose of bounding the coefficients of the vector-valued modular form $f$, it suffices to bound the asymptotic behaviour of scalar-valued modular forms for $\Gamma(N)$. There is the following standard bound available in the literature (cf. [Shimura1971, Lem. 3.62 p. 90], [ShimuraMF_HalfIntegralWeight, p. 447]).
@@
Also compare [Jin2018] deriving that if the component functions of a vector-valued modular form satisfy this bound, then it is a cusp form. Combining this bound with the bounds on coefficients of Eisenstein series yields the following result.
@@
In many cases there are sharper bounds on the asymptotic behaviour known, for instance, the bound $\mathcal{O}(n^{k/2-1/5})$ by Rankin in the even case (cf. [Rankin1939II, Thm. 2, p. 358]). However, even sharper bounds are to be found in the literature for forms to larger congruence subgroups. Consequently, a reduction to the setting of these bigger congruence subgroups is desirable.
Recall that $\Gamma(\operatorname{lev}(L))\trianglelefteq\Gamma_{1}(\operatorname{lev}(L))$ with abelian quotient, which remains true under the section $s$ from (2.5) so that we may, by a standard representation theoretic argument, reduce to the case of a modular form for $\Gamma_{1}(\operatorname{lev}(L))$ with finite character.
In a subsequent step, this information may be utilised to reduce to the even simpler case of $\Gamma_{0}(N)$ modular forms. However, first note that a scalar-valued modular form $f$ that has been transferred by the Petersson slash operator still possesses a Fourier expansion. Hence, it is meaningful to speak of a Fourier expansion of forms that have been altered by such a procedure. The proof idea of the following lemma is found in [Diamond2005, 1.2.11 p. 24].
@@
$f=\sum_{n\in\mathbb{N}/N}a(n)q^{n}\in\mathcal{S}_{k}(\Gamma_{1}(N),\psi).$
Assume there is some $l\in\mathbb{N}$ as well as some $\R\alpha\in\R$ such that for all characters $\chi:\Gamma_{0}(N^{2})\to\mathbb{T}$ that are trivial on $\Gamma_{1}(N^{2})$ and any choice $g=\sum_{n\in\mathbb{N}}b(n)q^{n}\in\mathcal{S}_{k}(\Gamma_{0}(N^{2}),\chi)$ we have $b(n^{l})\in\mathcal{O}(n^{\alpha})$. Then also
@@
    The inclusion follows from the condition on $N$ in conjunction with (2.5).
    Hence, for any choice of $\gamma\in s(\Gamma_{1}(N^{2}))$ we infer
@@
    Recall that by assumption $\psi$ was trivial on $s(\Gamma(N))$.
    As a result, we have that
    $f(N\tau)=f(M_{N}\tau)=\sum_{n\in\mathbb{N}/N}a(n)q^{Nn}\in\mathcal{S}_{k}(\Gamma_{1}(N^{2})).$
@@
Due to the structure of the Hecke algebra it suffices to find a similar bound for coefficients at primes that divide the level. In fact, by Atkin–Lehner theory, it suffices to do so for newforms. The following is due to Ogg and Li and part of [Li1974, Thm. 3 p. 295].
@@
However, for forms of half-integral weight, the situation is more delicate. Nevertheless, there are the following two bounds available in the literature. The first has been proven by Blomer and Harcos and is derived from [Blomer2008, Cor. 2 p. 55].

### v2

@@ 1. Introduction
Symmetric square $L$-series have already been constructed by Shimura [Shi73]. These were instrumental in establishing the Shimura lift – a pivotal correspondence between modular forms of half-integral and integral weights. This groundbreaking work catalysed further investigations, including Shimura’s own extensions [Shi75] and generalisations to higher rank groups in [BG92]. The present work addresses the natural extension of these $L$-series to the setting of vector-valued modular forms associated to the Weil representation. More explicitly, let $(L,\operatorname{q})$ denote a non-degenerate even lattice, $L^{\prime}$ denote its dual, $\mathfrak{e}_{\mu}$ a basis of the group algebra $\mathbb{C}[L^{\prime}/L]$, and $k\in\tfrac{1}{2}\mathbb{Z}$ a half-integral number. Further, if $\mathcal{M}_{L,k}$ represents the space of $\mathbb{C}[L^{\prime}/L]$ valued holomorphic modular forms of weight $k$ for the associated Weil representation, then any form $f\in\mathcal{M}_{L,k}$ has a Fourier expansion
@@
In extending symmetric square $L$-series from the scalar to the vector-valued setting, one might initially consider associating an $L$-series with each component function of $f$ individually, meaning for fixed $\mu\in L^{\prime}/L$ in (1.1). However, closer examination reveals that the natural analogues exhibit intertwining of the components. For clarity of exposition, we present a special instance of the series discussed in the main body of the paper. For an anisotropic vector $\ell\in L^{\prime}$ and $\overline{\ell}\in L^{\prime}/L$ its projection, we define
(1.2) $L_{(\overline{\ell},\operatorname{q}(\ell))}(f,s)=\sum_{n\in\mathbb{N}}\frac{a(n\overline{\ell},n^{2}\operatorname{q}(\ell))}{n^{s}}$
@@
A key feature of Shimura’s $L$-series is their admission of infinite product expansions when they are attached to Hecke eigenforms. We replicate this property in the vector-valued setting through careful analysis of Hecke operators, building upon the work of Bruinier and Stein (cf. [BS08, Ste15]). In the process, a bound on eigenvalues is derived (cf. Lemma 4.11). We state a simplified version of Theorem 5.12.
@@
$=\;$ $a(\lambda,\operatorname{q}(\ell))\cdot\prod_{p\nmid\operatorname{lev}(L)}\frac{1+\delta_{p\nmid\operatorname{q}(\ell)}\frac{G_{\mathcal{L}}(1)}{G_{\mathcal{L}}(p)}\cdot p^{k-1-s}}{1-\left(\frac{\sigma_{p}}{p^{k-1}}-(1-p^{-1})\frac{G_{\mathcal{L}}(1)}{G_{\mathcal{L}}(p)}\right)\cdot p^{k-1-s}+p^{2(k-1-s)}}$
@@ 2. Setting
We let $(L,\operatorname{q})$ denote a non-degenerate even quadratic lattice of signature $({m^{+}},{m^{-}})$ and denote its bilinear form by $\operatorname{b}$ such that $\operatorname{q}(x)\coloneq\operatorname{b}(x,x)/2$. The dual lattice $L^{\prime}\subset L\otimes_{\mathbb{Z}}\mathbb{Q}$ represents $\operatorname{Hom}_{\mathbb{Z}}(L,\mathbb{Z})$ via $\operatorname{b}$ and we write $\operatorname{lev}(L)$ for the level of the lattice, being the smallest positive integer such $(L^{\prime},\operatorname{lev}(L)\cdot\operatorname{q})$ is even. Recall the inclusion $L\leq L^{\prime}$ and note that $\mathcal{L}\coloneq L^{\prime}/L$ is a finite abelian group whose order is the absolute value of the determinant of a representing matrix of $\operatorname{b}$. Further, the quadratic form $\operatorname{q}$ descends to a quadratic form $\operatorname{\overline{q}}:\mathcal{L}\to\mathbb{Q}/\mathbb{Z}$ that is non-degenerate and such a pair $(\mathcal{L},\operatorname{\overline{q}})$ is called a discriminant form. In fact, any discriminant form arises from a non-degenerate even lattice in the above fashion. If we set $\operatorname{sig}(\mathcal{L})={m^{+}}-{m^{-}}\mod 8$ for some lattice $L$ inducing $\mathcal{L}$, this quantity is well defined by Milgram’s formula. For a natural number $n\in\mathbb{N}$ denote by $L(n)$ the pair $(L,n\cdot\operatorname{q})$ and set $\mathcal{L}(n)\coloneq L(n)^{\prime}/L(n)$.
@@
[…]
and find that $^{n}\!\mathcal{L}$} is exactly the orthogonal complement of $_{n}\mathcal{L}$}. We denote by $\mathbb{C}[\mathcal{L}]$ the group algebra of $\mathcal{L}$ with standard basis $\mathfrak{e}_{\lambda}$ for $\lambda\in\mathcal{L}$. There is a representation of $\operatorname{SL}_{2}(\mathbb{Z})$ on the group algebra $\mathbb{C}[\mathcal{L}]$ first described by Schoeneberg. In fact, to obtain a well-defined representation when $\operatorname{sig}(\mathcal{L})$ is odd, it must be considered on the metaplectic extension, rather than on $\operatorname{SL}_{2}$ directly. Let $\operatorname{\overline{GL_{2}^{+}}}(\mathbb{R})$ denote the metaplectic double cover of $\operatorname{GL}_{2}(\mathbb{R})^{+}$. A model is given by pairs $(\gamma,\phi)$, where $\gamma\in\operatorname{GL}_{2}(\mathbb{R})^{+}$ and $\phi:\mathbb{H}\to\mathbb{C}$ is a holomorphic square root of the factor of automorphy $j(\gamma,\tau)$, meaning $\phi^{2}(\tau)=j(\gamma,\tau)$. Multiplication is then declared by
@@
Projection to the first component $\operatorname{\overline{GL_{2}^{+}}}\ni(\gamma,\phi)\mapsto\gamma\in\operatorname{GL}_{2}^{+}(\mathbb{R})$ defines a covering. The action of an element $\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)=\gamma\in\operatorname{GL}_{2}(\mathbb{R})$ on the upper half plane $\mathbb{H}$ is given by Möbius transform $\tau\mapsto(a\tau+b)/(c\tau+d)$ with diagonal invariance group $\mathbb{R}^{\times}\operatorname{\hookrightarrow}\operatorname{GL}_{2}^{+}(\mathbb{R})$ and induces an action on $\operatorname{\overline{GL_{2}^{+}}}(\mathbb{R})$ via pullback. For the preimage of $\Gamma(1)\coloneq\operatorname{SL}_{2}(\mathbb{Z})$ in $\operatorname{\overline{GL_{2}^{+}}}(\mathbb{R})$, we write $\operatorname{Mp}_{2}(\mathbb{Z})$. This group is generated by the two elements $\overline{T}=\left(\left(\begin{smallmatrix}1&1\\ 0&1\end{smallmatrix}\right),1\right)$ and $\overline{S}=\left(\left(\begin{smallmatrix}0&-1\\ 1&0\end{smallmatrix}\right),\sqrt{\tau}\right)$, where $\sqrt{\tau}$ denotes the standard holomorphic square root. These elements project to the standard generators of $\operatorname{SL}_{2}(\mathbb{Z})$ and $\overline{S}^{2}=\overline{Z}=\left(-\left(\begin{smallmatrix}1&0\\ 0&1\end{smallmatrix}\right),i\right)$ generates the centre of $\operatorname{Mp}_{2}(\mathbb{Z})$.
@@
In case of odd rank, the oddity formula [Joh98, Chap. 15 p. 383 (30)] implies $4\mid N$, in particular, $\mathcal{L}$ contains $2$-adic Jordan components. In this case, there is a well known section
@@
with $\epsilon_{d}=1$ or $i$, depending on whether $d\equiv 1$ or $3\mod 4$. The argument at the end of [Bor00, Thm. 5.4 p. 330] yields that $\rho$$_{\mathcal{L}}$ is trivial on $s(\Gamma(N))$ and factors through the central extension of $\operatorname{SL}_{2}(\mathbb{Z}/N\mathbb{Z})$ by $\{\pm 1\}$ given by
@@
In fact, the section in (2.5) may be considered on $\Gamma_{0}(4)$, provided a larger central extension $\mathcal{G}$ of $\operatorname{SL}_{2}(\mathbb{R})$ is considered which is given in [Shi73, p. 444].
@@
Let $\gamma\in\operatorname{GL}_{2}^{+}(\mathbb{R})$ and $f:\mathbb{H}\to\mathbb{C}$ be a function. Define
@@
Finally, a holomorphic modular form $f:\mathbb{H}\to\mathbb{C}$ of weight $k$ for some congruence subgroup $\Gamma\leq\Gamma(1)$ with character $\chi:\Gamma\to\mathbb{T}$ is a holomorphic function such that for all $\gamma\in\Gamma$ we have $f|_{k}\gamma=\chi(\gamma)f$ and $f$ is holomorphic at the cusps of $\Gamma$. The space of such functions is denoted by $\mathcal{M}_{k}(\Gamma,\chi)$ and the subspace of elements vanishing at the cusps is denoted by $\mathcal{S}_{k}(\Gamma,\chi)$ and called the space of cusp forms. For the technical modifications required for forms of half-integral weight, we refer to [Shi73, p. 444]. Note that Definition 2.3 may also be generalised to the extension $\mathcal{G}$ of $\operatorname{GL}_{2}^{+}(\mathbb{R})$ that Shimura uses.
@@ 2.1. Gauss sums
This is a complementary section containing different Gauss sums appearing in section 4. The presented sums and some relations are already found in [Ste15, Sec. 4] including further computations which we will not repeat.
@@
Let $(L,\operatorname{q})$ be a quadratic $\mathbb{Z}$-lattice of rank $m\in\mathbb{N}$. Let $n\in\mathbb{N}$, $p$ be a prime, and consider $h\in\mathbb{Z}$ plus a Dirichlet Character $\chi$ modulo $n$. Define
@@
[…]
@@
$g_{p}\left[n,\chi_{1},h\right]$ $=\begin{cases}p^{n-1}(p-1),&\textnormal{ if }\nu_{p}(h)\geq n,\\ -p^{n-1},&\textnormal{ if }\nu_{p}(h)=n-1,\\ 0,&\textnormal{ if }\nu_{p}(h)<n-1.\end{cases}$
$g_{p}\left[n,\chi_{p},h\right]$ $=\begin{cases}p^{n-1/2}\left(\frac{h/p^{n-1}}{p}\right)\epsilon_{p},&\textnormal{ if }\nu_{p}(h)=n-1,\\ 0,&\textnormal{ if }\nu_{p}(h)\neq n-1.\end{cases}$
@@
Let $(L,\operatorname{q})$ be an integral $\mathbb{Z}$-lattice. By [Ste15, Lemma 4.3] for any $n\in\mathbb{N}$ and odd prime $p$, there is an orthogonal decomposition of $L/p^{n}L$ into $\mathbb{Z}/p^{n}\mathbb{Z}$ submodules
@@
where $L_{i}=(\mathbb{Z}/p^{n}\mathbb{Z})v_{i}$ is one dimensional with $\operatorname{b}(v_{i},v_{i})\in(\mathbb{Z}/p^{n}\mathbb{Z})^{\times}$, $M_{j}=(\mathbb{Z}/p^{n}\mathbb{Z})v_{j}$ with $\operatorname{b}(v_{j},v_{j})\in p^{k}(\mathbb{Z}/p^{n}\mathbb{Z})^{\times}$ where $1\leq k\leq n-1$, and $\operatorname{b}(N,N)\subseteq p^{n}\mathbb{Z}$. We may assume the $M_{j}$ being sorted with respect to increasing valuation $k$ of $\operatorname{q}(v_{j})$. Define
@@
In [BS08], the authors compare [McG03, Lem. 4.6 p. 115] to [Bor00, Thm. 5.4 p. 329] in order to conclude that for $\gcd(d,\operatorname{lev}(\mathcal{L}))=1$ the following association defines a character:
@@
Let $\mathcal{L}$ be the discriminant form of a lattice $L$ of rank $m$ and $d\in\mathbb{N}$. Then, by [Bar03, Thm. 5.2.2] we find
@@
Only the last equation has to be proven. By [Sch09, Prop. 3.8] we find the identity $|G_{\mathcal{L}}(d)|=\sqrt{|{{}_{d}\mathcal{L}}||\mathcal{L}|}$. As a consequence,
@@
Do note that explicit formulae for a more general version of $G_{\mathcal{L}}$ are also contained in [Sch09, Thm. 3.9 p. 11].
@@ 3. Modular forms
where $\overline{\mu}$ refers to the image of $\mu\in L^{\prime}/M$ under the projection to $L^{\prime}/L$ (cf. [Sch15, Sec. 4]). This construction plays a role in Subsection 5.1 and there are further means to construct induced vector-valued modular forms from existing ones described in [Sch15, Sec. 4]. However, a concrete class which is vital for our purposes is the following.
@@
(3.3) $\theta_{L,\lambda}(\tau)\coloneq\;$ $v^{\tfrac{{m^{-}}}{2}}\cdot\sum_{l\in L+\lambda}e\left(\tau\operatorname{q}(l)\right),\qquad\Theta_{L}\coloneq\sum_{\lambda\in L^{\prime}/L}\theta_{L,\lambda}(\tau)\mathfrak{e}_{\lambda}.$
These theta series $\Theta_{L}$ are modular forms of weight ${m^{+}}/2$ (cf. [Bor98, Thm. 4.1 p. 505]) and their Fourier coefficients equal representation numbers of the lattice $L$. Comparing these theta forms to Eisenstein series will yield formulae for the representation numbers in terms of divisor functions up to the coefficients of a cusp form. Thus, deriving asymptotic bounds on the Fourier coefficients of cusp forms is vital for quantifying these formulae. Additionally, these bounds serve, in our context, as a basis for proving convergence results for $L$-series associated to cusp forms. We will reduce this question to the scalar-valued case and begin with stating the following observation.
Proposition 3.3 ([Bor00, Thm. 5.4]).
@@
where $\chi_{\mathcal{L}}$ denotes the character of $\overline{\Gamma}_{0}(N)$ defined in [Bor00, Thm. 5.4].
In fact, in case of $4\nmid N$, the character $\chi_{\mathcal{L}}$ is trivial on the preimage of $\Gamma(N)$ in $\operatorname{Mp}_{2}(\mathbb{Z})$. In the other case, meaning $4\mid N$, the section $s(\Gamma(N))$ in $\operatorname{Mp}_{2}(\mathbb{Z})$ is contained in the kernel of the character (cf. [Bor00, Lem. 5.3 p. 329]). Hence, the component functions $f_{\lambda}$ of a vector-valued modular form $f\in\mathcal{M}_{L,k}$ are actually contained in $\mathcal{M}_{k}(\Gamma(\operatorname{lev}(L)))$.
As a consequence, for the purpose of asymptotically bounding the coefficients of the vector-valued modular form $f$, it suffices to bound the asymptotic behaviour of coefficients of scalar-valued modular forms for $\Gamma(N)$. There is the following standard bound available in the literature (cf. [Shi71, Lem. 3.62 p. 90], [Shi73, p. 447]).
@@
Also compare [JLL18] deriving that if the component functions of a vector-valued modular form satisfy this bound, then it is a cusp form. Combining this bound with the bounds on coefficients of Eisenstein series yields the following result.
@@
In many cases there are sharper bounds on the asymptotic behaviour known, for instance, the bound $\mathcal{O}(n^{k/2-1/5})$ by Rankin in the even case (cf. [Ran39, Thm. 2, p. 358]). However, even sharper bounds are to be found in the literature for forms to larger congruence subgroups. Consequently, a reduction to the setting of these bigger congruence subgroups is desirable. Recall that $\Gamma(\operatorname{lev}(L))\trianglelefteq\Gamma_{1}(\operatorname{lev}(L))$ with abelian quotient, which remains true under the section $s$ from (2.5) so that we may, by a standard representation theoretic argument, reduce to the case of a modular form for $\Gamma_{1}(\operatorname{lev}(L))$ with finite character.
In a subsequent step, this information may be utilised to reduce to the even simpler case of $\Gamma_{0}(N)$ modular forms. However, first note that a scalar-valued modular form $f$ that has been transferred by the Petersson slash operator still possesses a Fourier expansion. Hence, it is meaningful to speak of a Fourier expansion of forms that have been altered by such a procedure. The proof idea of the following lemma is found in [DS05, 1.2.11 p. 24].
@@
$f=\sum_{n\in\mathbb{N}/N}a(n)e(n\tau)\in\mathcal{S}_{k}(\Gamma_{1}(N),\psi).$
Assume there is some $l\in\mathbb{N}$ as well as some $\alpha\in\mathbb{R}$ such that for all characters $\chi:\Gamma_{0}(N^{2})\to\mathbb{T}$ that are trivial on $\Gamma_{1}(N^{2})$ and any choice $g=\sum_{n\in\mathbb{N}}b(n)q^{n}\in\mathcal{S}_{k}(\Gamma_{0}(N^{2}),\chi)$ we have $b(n^{l})\in\mathcal{O}(n^{\alpha})$. Then also
@@
The inclusion follows from the condition on $N$ in conjunction with (2.5). Hence, for any choice of $\gamma\in s(\Gamma_{1}(N^{2}))$ we infer
@@
Recall that by assumption $\psi$ was trivial on $s(\Gamma(N))$. As a result, we have that
$f(N\tau)=f(M_{N}\tau)=\sum_{n\in\mathbb{N}/N}a(n)e(Nn\tau)\in\mathcal{S}_{k}(\Gamma_{1}(N^{2})).$
@@
Due to the structure of the Hecke algebra it suffices to find a similar bound for coefficients at primes that divide the level. In fact, by Atkin–Lehner theory, it suffices to do so for newforms. The following is due to Ogg and Li and part of [Li74, Thm. 3 p. 295].
@@
However, for forms of half-integral weight, the situation is more delicate. Nevertheless, there are the following two bounds available in the literature. The first has been proven by Blomer and Harcos and is derived from [BH08, Cor. 2 p. 55].